# From Samples to Scenarios: A New Paradigm for Probabilistic Forecasting

### v1

@@ 1 Introduction
Probabilistic time series forecasting is fundamental to optimal decision-making under uncertainty, as it describes the likelihood of future outcomes (Gneiting & Katzfuss, 2014; Hyndman & Athanasopoulos, 2021). Although this problem has been studied extensively within the machine learning community, current approaches tend to rely on a predefined predictive distribution or sampling approximation (Kong et al., 2025; Fang & Wang, 2020; Lim & Zohren, 2021). These strategies have led to three main categories of models: (i) Parametric Distribution Models, assumes that the predictive distribution conforms to a predefined parametric family, such as a Gaussian (Salinas et al., 2020); (ii) Generative Models, such as diffusion-based models (Rasul et al., 2021), which learn an iterative process to generate samples from the latent distribution without explicitly defining its density function; and (iii) Structured Probabilistic Models, such as Flow-based Models (Rasul et al., 2020; Ashok et al., 2023), which learns a continuous probability density field, from which trajectories are then sampled.
However, the reliance on sampling introduces challenges (Cortés et al., 2025). While the alternative of a predefined distribution is not widely discussed by state-of-the-art methods for its evident inflexibility (Zhang et al., 2024; Ashok et al., 2023), the sampling paradigm suffers from three primary limitations, as shown in Figure 1: (i) Probability Absence. The most significant issue is that the generated trajectories are not paired with their probability of occurrence. Although confidence intervals can be inferred from a large set of samples, this process is indirect, computationally intensive, and lacks the intuitiveness of a direct scenario-probability mapping. (ii) Coverage Inadequacy. A finite set of samples may fail to represent low-probability, high-impact tail events. This is a critical failure for applications where preparing for rare occurrences is paramount, such as extreme weather or stock market volatility (Cortés et al., 2025). (iii) Inference Cost. The process of generating multiple samples is often expensive, with costs scaling to the number of samples required. This expense exacerbates the two preceding issues in practical applications, limiting the reliability and utility of such forecasts (Chen & Boccelli, 2018; Ashok et al., 2023).
@@ 2 Related Works
Generative Models represent the predictive distribution implicitly through a learned sampling process (Ho et al., 2020). For instance, Rasul et al. (2021) (TimeGrad) and Alcaraz & Strodthoff (2022) (SSSD) pioneered the use of conditional diffusion models for probabilistic forecasting and imputation. Recent advancements have focused on adapting the diffusion process to the sequential nature of time series (Gao et al., 2025; Biloš et al., 2023). Another direction focuses on enhancing the conditioning mechanism (Kollovieh et al., 2023; Liu et al., 2025) and adding cross-modal visual information (Ruan et al., 2025). Further adaptations include designing non-stationary processes (Ye et al., 2025) and specializing models for tasks like imputation, such as CSDI (Tashiro et al., 2021). Related work also employs Variational Autoencoders (VAEs), such as GP-VAE (Fortuin et al., 2020). While powerful, all these methods produce forecasts via an iterative sampling procedure and do not provide an explicit probability density for any given trajectory.
Structured Probabilistic Models are a class of methods that learn an explicit, continuous probability density field over the forecast horizon, primarily including flow-based models and copula-based models. Flow-based models learn a distribution transformation (Papamakarios et al., 2021), with recent applications in time series forecasting, such as TempFlow, using techniques like conditioned normalizing flows and flow matching (Rasul et al., 2020; Kollovieh et al., 2024). Copula-based models, which have a long history in econometrics and finance (Patton, 2012; Größer & Okhrin, 2022), construct a joint distribution by a copula (Salmon & Bouyé, 2008; Bouyé et al., 2008; Wang & Tao, 2020). With neural networks involved to model the copula (Wen & Torkkola, 2019; Krupskii & Joe, 2020; Mayer & Wied, 2023; Toubeau et al., 2019), recent research leads to fully neural, non-parametric approaches like TACTiS (Drouin et al., 2022) and its successor, TACTiS-2 (Ashok et al., 2023). Despite their different formulations, models in this category still rely on a sampling procedure, drawing from a continuous probability density field to obtain future trajectories.
@@ 3.3 Unified Evaluation Framework
To quantitatively measure the limitations of sampling-based methods, we establish an evaluation framework by adapting two complementary metrics. We use the standard metric for overall forecast quality (Zhang et al., 2024; Zheng & Sun, 2025), the CRPS, to assess Probability Absence. Concurrently, we use Distortion, defined as the error of the best single trajectory in a set, to assess Coverage Inadequacy (Cortés et al., 2025). For both metrics, we provide distinct but comparable formulations for the sampling-based and Probabilistic Scenarios paradigms.
@@ 5.1 Basic Setup
Data. Following the recent benchmark for probabilistic time series forecasting, ProTS (Zhang et al., 2024), we evaluate our model on five datasets: Electricity (Elec.), Exchange (Exch.), Solar(Sol.), Traffic(Traf.), and Wikipedia (Wiki.). These are benchmark datasets taken from the GluonTS library (Alexandrov et al., 2020), preprocessed exactly as in prior works (Gasthaus et al., 2019). A detailed analysis of dataset properties cited from previous work is provided in the Appendix C.1. Following prior work, we set the forecast horizon to 24 (hours) for the hourly datasets (Electricity, Solar, Traffic) and 30 (days) for the daily datasets (Exchange, Wikipedia) (Zhang et al., 2024). For TimePrism, the input length is set equal to the forecast horizon. Other baselines may use longer context lengths as lagged features, for which we adhere to their configurations in prior work (Cortés et al., 2025). Notably, TimePrism achieves strong performance even with less information input. For a comprehensive comparison, we also provide results in the Appendix D.2.1, where TimePrism uses an input length comparable to that of the baselines.
@@
Baselines. For a comprehensive comparison, we select seven models covering all three categories discussed in our related work. ETS (Hyndman et al., 2008) serves as a non-neural baseline. DeepAR (Salinas et al., 2020) represents parametric distribution models. TimeGrad (Rasul et al., 2021) is a diffusion-based generative model. TempFlow (Rasul et al., 2020), Transformer TempFlow (Trf.Flow), and TACTiS-2 (Ashok et al., 2023) are structured probabilistic models. Tempflow is implemented with Long Short-Term Memory (LSTM) backbone (Hochreiter & Schmidhuber, 1997) and Trf.Flow is implemented with a Transformer backbone (Vaswani et al., 2017), as in Rasul et al. (2020). Finally, TimeMCL (Cortés et al., 2025) represents multi-choice learning models.
Training Details. All models are trained using the Adam optimizer with an initial learning rate of $10^{-3}$ for 200 epochs. Given its lack of hidden layers, the number of scenarios $N$ is the primary tunable hyperparameter for TimePrism. In this section, TimePrism uses $N=625$ scenarios, composed of $M=25$ trend and $K=25$ seasonal components. In practice, if the number of distinct future scenarios is known a priori, $N$ can be set to match this number; otherwise, as in our benchmark datasets, $N$ should be set to a value large enough to allow the model to learn on its own. Further training details and analysis on $N$ are included in the Appendix C.3.
@@
5.2 Main Results
Coverage Inadequacy and Distortion. To assess Coverage Inadequacy, we use the Distortion metric, with results presented in Table 2. TimePrism achieves the state-of-the-art result across all five datasets, demonstrating its superior ability to generate a diverse set of scenarios that covers the ground truth. This is because our reconstruction loss, $\mathcal{L}_{\text{recon}}$, allows the model not to be heavily penalized for predicting a plausible but non-realized future, in datasets containing similar histories but diverse futures. Notably, TimeMCL, based on discrete scenarios but without explicit probabilities, achieves the second-best result on three datasets in Distortion. This suggests that discrete-hypothesis models have an inherent advantage in forecast coverage over purely sampling-based models.
@@ 6 Conclusions and Discussion
Conclusion. Probabilistic time series forecasting is crucial for reliable decision-making. While powerful, current SOTA methods predominantly rely on sampling, a paradigm that faces limitations of Probability Absence, Coverage Inadequacy, and Inference Cost. To address these challenges, we introduced the Probabilistic Scenarios paradigm. This paradigm operates by directly producing a set of {Scenario, Probability} pairs in a single forward pass, without reliance on sampling. We validated this new paradigm with TimePrism, a simple linear model. Evaluated under our unified framework, TimePrism addresses these challenges and demonstrates the potential of the new paradigm.
Discussion. We conclude by discussing the applicability of our approach and outlining future work:
  • Reason of Effectiveness. The strong performance of TimePrism stems from the paradigm’s reframing of the learning objective. Instead of learning to model an entire continuous probability space, the model is learning a more structured problem: a probability distribution over a discrete set of scenarios. This shift reduces the required model capacity, allowing a simple linear architecture to achieve strong results.
  • Applicability of the TimePrism. The intentionally simple structure of TimePrism, while effective for validating the paradigm, may have limitations in more complex scenarios, such as those with extremely high dimensionality, or series lacking trend or seasonal patterns.
  • Future Models within the new Paradigm. TimePrism serves only as a proof-of-concept. The true potential of Probabilistic Scenarios lies in its application to more powerful backbones. Future work could integrate this paradigm with state-of-the-art architectures like Transformers, Diffusion, or Flow Matching models to unlock new levels of multivariate performance.
  • Future Refinements of the new Paradigm. The paradigm itself can be further enhanced. For instance, developing methods to adaptively determine the number of scenarios based on data complexity could improve its practical utility.
In summary, our work provides a practical alternative to sampling and broadens the conceptual landscape of probabilistic forecasting, establishing a promising foundation for future research.
@@ A.2.1 Optimal Scenarios via Reconstruction Loss
Proposition 1. Under standard stationarity assumptions, the optimal set of $M$ trend forecasts $\mathcal{M}=\{\mathbf{y}_{\text{t},m}\}_{m=1}^{M}$ and $K$ seasonal forecasts $\mathcal{K}=\{\mathbf{y}_{\text{s},k}\}_{k=1}^{K}$ are those whose $N=M\cdot K$ combinations form a centroidal Voronoi tessellation of the space of future trajectories, conditioned on the input history. Each combined scenario $\mathbf{y}_{\text{t},m}+\mathbf{y}_{\text{s},k}$ converges to the conditional mean of its corresponding Voronoi region.
@@
The min operator partitions the space of future trajectories, for a given $\mathbf{x}$, into $N$ Voronoi regions, $\{R_{n}(\mathbf{x})\}_{n=1}^{N}$, where each region $R_{n}(\mathbf{x})$ contains all trajectories $\mathbf{y}_{\text{gt}}$ for which the $n$-th scenario is the winner. The inner expectation can then be rewritten as a sum of integrals over these regions:
$\mathbb{E}_{\mathbf{y}_{\text{gt}}|\mathbf{x}}[\dots]=\sum_{n=1}^{N}\int_{R_{n}(\mathbf{x})}\left\|\mathbf{y}_{\text{gt}}-\mathbf{y}_{n}(\mathbf{x})\right\|_{2}^{2}p(\mathbf{y}_{\text{gt}}|\mathbf{x})d\mathbf{y}_{\text{gt}}$ (13)
@@
This demonstrates that for any given history $\mathbf{x}$, the optimal scenarios generated by our model must be the conditional means of their respective Voronoi regions (Du et al., 1999). In geometric terms, the set of $N$ scenarios acts as a set of centers that partition the high-dimensional space of all possible futures into $N$ distinct regions, known as a Voronoi tessellation. Each region consists of all future trajectories that are closer to one specific scenario than to any other. Our result shows that the WTA training objective effectively drives the model to find an optimal set of "cluster centers" (our scenarios) that best represent the underlying structure of the data, where "best" is defined in the sense of minimizing the expected squared error, akin to the objective in k-means clustering (Cortés et al., 2025; Arthur & Vassilvitskii, 2007).

### v2

@@ 1 Introduction
Probabilistic time series forecasting is fundamental to optimal decision-making under uncertainty, as it describes the likelihood of future outcomes (Gneiting and Katzfuss, 2014; Hyndman and Athanasopoulos, 2021). Although this problem has been studied extensively within the machine learning community, current approaches tend to rely on a predefined predictive distribution or sampling approximation (Kong et al., 2025; Fang and Wang, 2020; Lim and Zohren, 2021). These strategies have led to three main categories of models: (i) Parametric Distribution Models, assumes that the predictive distribution conforms to a predefined parametric family, such as a Gaussian (Salinas et al., 2020); (ii) Generative Models, such as diffusion-based models (Rasul et al., 2021), which learn an iterative process to generate samples from the latent distribution without explicitly defining its density function; and (iii) Structured Probabilistic Models, such as Flow-based Models (Rasul et al., 2020; Ashok et al., 2023), which learns a continuous probability density field, from which trajectories are then sampled.
However, the reliance on sampling introduces challenges (Cortés et al., 2025). While the alternative of a predefined distribution is not widely discussed by state-of-the-art methods for its evident inflexibility (Zhang et al., 2024; Ashok et al., 2023), the sampling paradigm suffers from three primary limitations, as shown in Figure 1: (i) Probability Absence. The most significant issue is that the generated trajectories are not paired with their probability of occurrence. Although confidence intervals can be inferred from a large set of samples, this process is indirect, computationally intensive, and lacks the intuitiveness of a direct scenario-probability mapping. (ii) Coverage Inadequacy. A finite set of samples may fail to represent low-probability, high-impact tail events. This is a critical failure for applications where preparing for rare occurrences is paramount, such as extreme weather or stock market volatility (Cortés et al., 2025). (iii) Inference Cost. The process of generating multiple samples is often expensive, with costs scaling to the number of samples required. This expense exacerbates the two preceding issues in practical applications, limiting the reliability and utility of such forecasts (Chen and Boccelli, 2018; Ashok et al., 2023).
@@ 2 Related Works
Generative Models represent the predictive distribution implicitly through a learned sampling process (Ho et al., 2020). For instance, Rasul et al. (2021) (TimeGrad) and Alcaraz and Strodthoff (2022) (SSSD) pioneered the use of conditional diffusion models for probabilistic forecasting and imputation. Recent advancements have focused on adapting the diffusion process to the sequential nature of time series (Gao et al., 2025; Biloš et al., 2023). Another direction focuses on enhancing the conditioning mechanism (Kollovieh et al., 2023; Liu et al., 2025) and adding cross-modal visual information (Ruan et al., 2025). Further adaptations include designing non-stationary processes (Ye et al., 2025) and specializing models for tasks like imputation, such as CSDI (Tashiro et al., 2021). Related work also employs Variational Autoencoders (VAEs), such as GP-VAE (Fortuin et al., 2020). While powerful, all these methods produce forecasts via an iterative sampling procedure and do not provide an explicit probability density for any given trajectory.
Structured Probabilistic Models are a class of methods that learn an explicit, continuous probability density field over the forecast horizon, primarily including flow-based models and copula-based models. Flow-based models learn a distribution transformation (Papamakarios et al., 2021), with recent applications in time series forecasting, such as TempFlow, using techniques like conditioned normalizing flows and flow matching (Rasul et al., 2020; Kollovieh et al., 2024). Copula-based models, which have a long history in econometrics and finance (Patton, 2012; Größer and Okhrin, 2022), construct a joint distribution by a copula (Salmon and Bouyé, 2008; Bouyé et al., 2008; Wang and Tao, 2020). With neural networks involved to model the copula (Wen and Torkkola, 2019; Krupskii and Joe, 2020; Mayer and Wied, 2023; Toubeau et al., 2019), recent research leads to fully neural, non-parametric approaches like TACTiS (Drouin et al., 2022) and its successor, TACTiS-2 (Ashok et al., 2023). Despite their different formulations, models in this category still rely on a sampling procedure, drawing from a continuous probability density field to obtain future trajectories.
@@ 3.3 Unified Evaluation Framework
To quantitatively measure the limitations of sampling-based methods, we establish an evaluation framework by adapting two complementary metrics. We use the standard metric for overall forecast quality (Zhang et al., 2024; Zheng and Sun, 2025), the CRPS, to assess Probability Absence. Concurrently, we use Distortion, defined as the error of the best single trajectory in a set, to assess Coverage Inadequacy (Cortés et al., 2025). For both metrics, we provide distinct but comparable formulations for the sampling-based and Probabilistic Scenarios paradigms.
@@ 5.1 Basic Setup
Data. Following the recent benchmark for probabilistic time series forecasting, ProTS (Zhang et al., 2024), we evaluate our model on five datasets: Electricity (Elec.), Exchange (Exch.), Solar(Sol.), Traffic(Traf.), and Wikipedia (Wiki.). These are benchmark datasets taken from the GluonTS library (Alexandrov et al., 2020), preprocessed exactly as in prior works (Gasthaus et al., 2019). A detailed analysis of dataset properties cited from previous work is provided in the Appendix C.1. Following prior work, we set the forecast horizon to 24 (hours) for the hourly datasets (Electricity, Solar, Traffic) and 30 (days) for the daily datasets (Exchange, Wikipedia) (Zhang et al., 2024). For TimePrism, the input length is set equal to the forecast horizon. Other baselines may use longer context lengths as lagged features, for which we adhere to their configurations in prior work (Cortés et al., 2025). Notably, TimePrism achieves strong performance even with less information input. For a comprehensive comparison, we also provide results in the Appendix D.2, where TimePrism uses an input length comparable to that of the baselines.
@@
Baselines. For a comprehensive comparison, we select seven models covering all three categories discussed in our related work. ETS (Hyndman et al., 2008) serves as a non-neural baseline. DeepAR (Salinas et al., 2020) represents parametric distribution models. TimeGrad (Rasul et al., 2021) is a diffusion-based generative model. TempFlow (Rasul et al., 2020), Transformer TempFlow (Trf.Flow), and TACTiS-2 (Ashok et al., 2023) are structured probabilistic models. Tempflow is implemented with Long Short-Term Memory (LSTM) backbone (Hochreiter and Schmidhuber, 1997) and Trf.Flow is implemented with a Transformer backbone (Vaswani et al., 2017), as in Rasul et al. (2020). Finally, TimeMCL (Cortés et al., 2025) represents multi-choice learning models.
Training Details. All models are trained using the Adam optimizer with an initial learning rate of $10^{-3}$ for 200 epochs. Given its lack of hidden layers, the number of scenarios $N$ is the primary tunable hyperparameter for TimePrism. In this section, TimePrism uses $N=625$ scenarios, composed of $M=25$ trend and $K=25$ seasonal components. In practice, if the number of distinct future scenarios is known a priori, $N$ can be set to match this number; otherwise, as in our benchmark datasets, $N$ should be set to a value large enough to allow the model to learn on its own. Further training details and analysis on $N$ are included in the Appendix C.3.To ensure a fair comparison, considering that TimeMCL employs 16 hypotheses in its original implementation, we specifically include a variant with $N=16$, denoted as TimePrism-16. This serves to validate the effectiveness of the new paradigm, demonstrating that it functions effectively with a simple structure and without requiring a large number of parameters.
5.2 Main Results
@@
Coverage Inadequacy and Distortion. To assess Coverage Inadequacy, we use the Distortion metric, with results presented in Table 2. TimePrism achieves the state-of-the-art result across all five datasets, demonstrating its superior ability to generate a diverse set of scenarios that covers the ground truth. This is because our reconstruction loss, $\mathcal{L}_{\text{recon}}$, allows the model not to be heavily penalized for predicting a plausible but non-realized future, in datasets containing similar histories but diverse futures.
@@
5.3 Impact of Scenario Set Size
Table 4: Impact of Scenario Count (NN) on Performance and Complexity. This table presents a systematic ablation study across Electricity, Exchange, and Solar datasets, illustrating how model complexity and forecasting error (CRPS, Distortion) scale with the number of scenarios $N$.
N FLOPs Solar Electricity Exchange
CRPS Distortion CRPS Distortion CRPS Distortion
1 1.0x 0.199 $\pm$ 0.00 0.266 $\pm$ 0.00 0.409 $\pm$ 0.01 0.733 $\pm$ 0.03 0.596 $\pm$ 0.00 0.803 $\pm$ 0.00
16 4.2x 0.137 $\pm$ 0.00 0.307 $\pm$ 0.03 0.414 $\pm$ 0.10 0.911 $\pm$ 0.22 0.611 $\pm$ 0.04 0.920 $\pm$ 0.04
256 19.9x 0.0927 $\pm$ 0.00 0.158 $\pm$ 0.01 0.162 $\pm$ 0.01 0.267 $\pm$ 0.03 0.486 $\pm$ 0.00 0.666 $\pm$ 0.02
625 34.8x 0.0852 $\pm$ 0.00 0.101 $\pm$ 0.03 0.133 $\pm$ 0.01 0.211 $\pm$ 0.03 0.468 $\pm$ 0.01 0.595 $\pm$ 0.01
1024 48.3x 0.0822 $\pm$ 0.00 0.0917 $\pm$ 0.01 0.139 $\pm$ 0.01 0.212 $\pm$ 0.02 0.452 $\pm$ 0.00 0.583 $\pm$ 0.02
We conduct a systematic analysis to investigate the trade-off between the scenario set size $N$ and model performance. Table 4 summarizes the results across three representative datasets (Electricity, Exchange, and Solar) with $N$ ranging from 1 to 1024.
Complexity Scaling. A key advantage of our combinatorial architecture ($N=M\times K$) is its efficiency. The parameter complexity of the shared basis layers (Trend and Season) scales as $\mathcal{O}(\sqrt{N})$, while only the probability head scales linearly as $\mathcal{O}(N)$. Consequently, the overall model complexity grows favorably between $\mathcal{O}(N^{1/2})$ and $\mathcal{O}(N)$, allowing for large scenario sets.
Performance Trends. Increasing $N$ generally leads to lower CRPS and Distortion errors, as a larger discrete set can approximate the continuous probability space with higher fidelity. However, we observe diminishing returns: the performance gains tend to plateau around $N=625$. Beyond this point, the marginal benefit of adding scenarios decreases while the computational cost continues to rise. Based on this equilibrium, we adopted $N=625$ as the unified setting for our main experiments.
Dataset Dependence. The results also indicate that the "saturation point" varies slightly by dataset. For instance, the Solar dataset benefits more from a larger $N$ compared to the Exchange dataset. This suggests that the optimal $N$ is determined by the intrinsic complexity of the data, highlighting the potential for future work on adaptive mechanisms that dynamically adjust $N$.
@@ 6 Conclusions and Discussion
6.1 Discussion
Reason of Effectiveness. The strong performance of TimePrism stems from the paradigm’s reframing of the learning objective. Instead of learning to model an entire continuous probability space, the model is learning a more structured problem: a probability distribution over a discrete set of scenarios. This concept parallels Vector Quantization (VQ) techniques in representation learning, most notably VQ-VAE (van den Oord et al., 2017), but applies the discretization directly to the output trajectory space rather than a latent space (see Appendix A.3 for a detailed discussion). This shift reduces the required model capacity, allowing a simple linear architecture to achieve strong results.
Limitations:
  • Dataset Applicability. The intentionally simple structure of TimePrism, while effective for validating the paradigm, may have limitations in more complex scenarios, such as those with extremely high dimensionality, or series lacking trend or seasonal patterns.
  • Structural Rigidity. As a linear model, the current version of TimePrism requires fixed-length inputs and prediction horizons, limiting its flexibility in scenarios where variable-length contexts are available during inference.
  • Simplified Multivariate Modeling. Our current implementation utilizes a weight-sharing strategy. We believe there is significant room for improvement by incorporating more sophisticated channel-mixing mechanisms to model cross-variate relationships.
Future Works:
  • Models within the new Paradigm. TimePrism serves only as a proof-of-concept. The true potential of Probabilistic Scenarios lies in its application to more powerful backbones. Future work could integrate this paradigm with state-of-the-art architectures like Transformers, Diffusion, or Flow Matching models to unlock new levels of multivariate performance.
  • Refinements of the new Paradigm. The paradigm itself can be further enhanced. For instance, developing methods to adaptively determine the number of scenarios based on data complexity could improve its practical utility.
  • Decision-Centric Assessment. Metrics like CRPS and Distortion may not fully reflect the downstream utility of probabilistic forecasts in real-world environments. In future work, decision-centric metrics can be incorporated, such as tail-risk assessment and utility-based scores. Furthermore, we plan to explore the direct integration of our Probabilistic Scenarios paradigm into real-world decision-making to demonstrate its practical value beyond pure forecasting accuracy.
6.2 Conclusion
Probabilistic time series forecasting is crucial for reliable decision-making. While powerful, current SOTA methods predominantly rely on sampling, a paradigm that faces limitations of Probability Absence, Coverage Inadequacy, and Inference Cost. To address these challenges, we introduced the Probabilistic Scenarios paradigm. This paradigm operates by directly producing a set of {Scenario, Probability} pairs in a single forward pass, without reliance on sampling. We validated this new paradigm with TimePrism, a simple linear model. Evaluated under our unified framework, TimePrism addresses these challenges and demonstrates the potential of the new paradigm. In summary, our work provides a practical alternative to sampling and broadens the conceptual landscape of probabilistic forecasting, establishing a promising foundation for future research.
@@ 6.2 Conclusion
The code of this work is available at: https://github.com/Fifthky/TimePrism.
@@ A.2.1 Optimal Scenarios via Reconstruction Loss
Proposition 1. Assuming that the model parameters reach a local minimum of the reconstruction loss, a necessary condition is that the set of $N=M\cdot K$ combined scenarios forms a Centroidal Voronoi Tessellation (CVT) of the space of future trajectories, conditioned on the input history. Specifically, each combined scenario $\mathbf{y}_{\text{t},m}+\mathbf{y}_{\text{s},k}$ converges to the conditional mean of its corresponding Voronoi region.
@@
The min operator partitions the space of future trajectories, for a given $\mathbf{x}$, into $N$ Voronoi regions, $\{R_{n}(\mathbf{x})\}_{n=1}^{N}$. We formally define the Voronoi region for the $n$-th scenario as $R_{n}(\mathbf{x})=\{\mathbf{y}\in\mathbb{R}^{T\times D}\mid\|\mathbf{y}-\mathbf{y}_{n}(\mathbf{x})\|_{2}\leq\|\mathbf{y}-\mathbf{y}_{j}(\mathbf{x})\|_{2},\forall j\}$. Each region $R_{n}(\mathbf{x})$ contains all trajectories $\mathbf{y}_{\text{gt}}$ for which the $n$-th scenario is the winner. The inner expectation can then be rewritten as a sum of integrals over these regions:
$\mathbb{E}_{\mathbf{y}_{\text{gt}}|\mathbf{x}}\left[\min_{n=1\dots N}\left\|\mathbf{y}_{\text{gt}}-\mathbf{y}_{n}(\mathbf{x})\right\|_{2}^{2}\right]=\sum_{n=1}^{N}\int_{R_{n}(\mathbf{x})}\left\|\mathbf{y}_{\text{gt}}-\mathbf{y}_{n}(\mathbf{x})\right\|_{2}^{2}p(\mathbf{y}_{\text{gt}}|\mathbf{x})d\mathbf{y}_{\text{gt}}$ (13)
@@
This derivation holds provided that the Voronoi region has non-zero probability mass, that is, when $\int_{R_{n}(\mathbf{x})}p(\mathbf{y}_{\text{gt}}|\mathbf{x})d\mathbf{y}_{\text{gt}}\neq 0$. This demonstrates that for any given history $\mathbf{x}$, the scenarios generated by an optimal model must be the conditional means of their respective Voronoi regions (Du et al., 1999). In geometric terms, the set of $N$ scenarios acts as a set of centers that partition the high-dimensional space of all possible futures into $N$ distinct regions, known as a Voronoi tessellation. Each region consists of all future trajectories that are closer to one specific scenario than to any other. Our result shows that the WTA training objective effectively drives the model to find an optimal set of "cluster centers" (our scenarios) that best represent the underlying structure of the data, where "best" is defined in the sense of minimizing the expected squared error, akin to the objective in k-means clustering (Cortés et al., 2025; Arthur and Vassilvitskii, 2007).